# Support Recovery and $\ell_{2}$-Error Bound for Sparse Regression with Quadratic Measurements via Weakly-Convex-Concave Regularization

### v1

@@ 1 Introduction
We consider the following quadratic measurement regression model
@@
where $y_{i}\in\mathbb{R}$ denotes the observed response, $\bm{Z}_{i}\in\mathbb{R}^{d\times d}$ is a symmetric design matrix, $\bm{\beta}\in\mathbb{R}^{d}$ is the vector of unknown parameters representing true signals, and $\varepsilon_{i}\in\mathbb{R}$ is the random noise.
Model (1) arises in numerous applications in physics, engineering, and data science, including phase retrieval (Candes et al. 2015), generalized phase retrieval (Wang and Xu 2019), the unassigned distance geometry problem (Huang and Dokmanić 2021), and power system state estimation (Wang et al. 2019). Recent years have witnessed a surge of interests and methodological developments for this model. Notable advances include Thaker et al. (2020), Huang et al. (2020), Chen and Ng (2022), and Fan et al. (2025a), who proposed various algorithms and theoretical guarantees of optimal solutions under various problem settings. It is worth noting that when the design matrix can be decomposed as a outer product of a real vector $\bm{Z}_{i}=\bm{z}_{i}\bm{z}_{i}^{\top}$, model (1) is referred to as real-valued phase retrieval. In this case, it also represents a special case of the single index model, with index $\bm{z}_{i}^{\top}\bm{\beta}$ and square link function. Single index models are widely used in statistics and econometrics (Xu et al. 2022, Yang et al. 2024).
In high-dimensional regime where $d$ is large or even exceeds $n$, a common assumption is that the true signal is sparse, i.e., only a small number of its entries are non-zero. This sparsity assumption is crucial for ensuring identifiability and statistical efficiency, leading to the problem of sparse regression with quadratic measurements. Several studies have explored this challenging problem. For instance, Fan et al. (2018) studied an $\ell_{q}(0<q<1)$-regularized least squares method, established a weak oracle property and proposed a fixed-point iterative algorithm. Bolte et al. (2018) called the regularized least squares problem for model (1) quadratic inverse problem and proposed a Bregman proximal gradient algorithm (PGA). Subsequently, Zhang et al. (2023) and Ding et al. (2025) developed Bregman PGA to solve regularized problems including quadratic inverse problem as a special case. While existing literature has primarily focused on algorithmic development for the regularized least squares formulation of quadratic inverse problem (or model (1)), the corresponding statistical guarantees remain relatively underexplored. More recently, Chen et al. (2025) employed the thresholded Wirtinger flow algorithm originally proposed by (Cai et al. 2016) for sparse phase retrieval to solve the regularized least squares problem for model (1). A special case of sparse quadratic model is sparse phase retrieval that attracted considerable attention (Cai et al. 2016, 2022, Soltanolkotabi 2019, Xia and Xu 2021, Huang and Xu 2024).
To address sparse regression in linear models, various other regularization techniques have been developed, such as the $\ell_{q}(0<q<1)$ penalty (Frank and Friedman 1993), the SCAD (Fan and Li 2001), and the MCP (Zhang 2010). It is well-known that LASSO has constant derivative that induces persistent shrinkage and biased estimation(Zou 2006). In contrast, concave penalties (e.g., SCAD, MCP, and $\ell_{q}(0<q<1)$) have derivatives that vanish for large coefficients, thereby mitigating shrinkage bias and yielding nearly unbiased estimates. However, nonconvex penalties introduces computational challenges. In particular, highly nonconvex penalties may cause numerous local minima, and the convergence of first-order algorithm (e.g., PGA) relies on favorable properties of the objective function. One such property is weak convexity, which is a key condition that generally ensures the proximal operator is single-valued (Wang 2010, Khanh et al. 2025), under standard technical assumptions. If this condition is violated, the proximal operator may be multi-valued, which poses a potential risk that iterative algorithms could fail to converge to a meaningful solution.
This challenge motivates the study of regularizers that balance statistical and computational considerations. Loh and Wainwright (2015) developed a general statistical analysis for a class of weakly convex regularized M-estimators and established finite sample error bounds between any stationary point of the penalized estimator and the population-optimal solution without regularization (i.e., the solution minimizing the expected population risk). While their framework applies to a wide range of models, including linear and generalized linear models and graphical LASSO, it cannot be directly applied to sparse quadratic measurement problems. The main difficulty is that the least squares loss for this model leads to a highly nonconvex quartic polynomial that are computationally very challenging to minimize. In fact, as pointed out by Candes et al. (2015), it is already known that determining whether a stationary point of a quartic polynomial is a local minimizer is NP-hard. Moreover, Loh and Wainwright (2015) imposed an additional constraint, which introduces an extra tuning parameter that must be chosen carefully to ensure that the true value $\bm{\beta}^{*}$ is a feasible point, thereby further increasing the difficulty of solving the problem. Weakly convex regularization methods have also been widely used in sparse signal recovery (Yang et al. 2019, Komuro et al. 2022) and linear inverse problems (Shumaylov et al. 2024, Goujon et al. 2024, Ebner et al. 2025), but these works do not explicitly exploit concave penalties.
To mitigate the above mentioned numerical issues, in this paper, we propose a specific class of concave penalties that also satisfy weak convexity, termed Weakly-Convex-Concave Penalties (WCCP). The major advantage of this class is that it provides a flexible framework for analyzing local minimizers in sparse quadratic measurement problems, combining the statistical benefits of concavity with the algorithmic guarantees of weak convexity.
Specifically, in this paper, we study the regularization problem
@@
where $\lambda_{n}>0$ is a tuning parameter that encourages sparsity, and $P_{\lambda_{n}}(\cdot)$ is a WCCP. As mentioned before, model (1) constitutes a special case of the single index model with a known square link function. Therefore, the proposed method (2) has potential applications for single index models.
Our contributions are threefold. First, we bridge a notable gap in the literature by providing the first systematic statistical analysis of the WCCP-regularized estimator for the quadratic measurement model (1). We establish rigorous support recovery guarantees and $\ell_{2}$-error bounds for its local minimizer. we derive the statistical properties of the WCCP-regularized estimator for the quadratic measurement model (1) and establish support recovery and tight $\ell_{2}$-error bounds for the local minimizer. Second, we develop iterative algorithms tailored to different regularization functions that provide convergence guarantees. Third, we demonstrate the effectiveness of our method, particularly in the context of sparse phase retrieval, and show that it performs comparable or better than the existing LASSO-based approaches.
@@ 2 Finite Sample Statistical Results
In this section, we derive the finite sample properties and consistency of the weakly convex-concave estimator in the high-dimensional case where $d$ is large than $n$. We assume that $\ln d=o(n^{\alpha})$ with some constant $\alpha\in(0,1)$ and denote the number of non-zero elements of the true signal by $s$. Following literature (Huang et al. 2008, Fan et al. 2018), we assume that there exists constants $0<c_{1}\leq c_{2}<\infty$ such that
@@
Further regarding the observed response and design matrix, following Fan et al. (2018), the data are assumed to be standardized as
@@
The regularization function is coordinate separable:
@@
for some scalar function $p_{\lambda_{n}}$ which satisfies
@@
    The function $p_{\lambda_{n}}$ is concave on $(0,\infty)$ and there exists $\mu>0$ such that $p_{\lambda_{n}}(t)+\frac{\mu}{2}t^{2}$ is convex.
@@
    The function $p_{\lambda_{n}}$ satisfies $p_{\lambda_{n}}(0)=0$ and $p_{\lambda_{n}}(t)>0$ for $t>0$.
@@
    For $t>0$, the function $p_{\lambda_{n}}$ is non-decreasing and $t\mapsto\frac{p_{\lambda_{n}}(t)}{t}$ is non-increasing in $t$.
@@
    The function $p_{\lambda_{n}}$ is differentiable with derivative $p^{\prime}_{\lambda_{n}}$ for all $t>0$, and subdifferentiable at $t=0$, satisfying $\lim_{t\to 0^{+}}p^{\prime}_{\lambda_{n}}(t)=\lambda_{n}\varrho$ with $\varrho>0$. The function $p^{\prime}_{\lambda_{n}}$ is locally Lipschitz continuous in $(0,\infty)$.
Compared to the conditions of (Loh and Wainwright 2017, Loh 2017), Assumption 1 additionally requires the penalty function to be concave and its derivative $p_{\lambda_{n}}^{{}^{\prime}}$ to be locally Lipschitz continuous on the positive half-line. In fact, the condition used by (Loh and Wainwright 2017, Loh 2017) is relaxed enough to include the LASSO penalty. As mentioned earlier, concave penalties are known to possess stronger variable selection capabilities. Therefore, we explicitly emphasize the concavity of the penalty function in our framework. Compared to a general concave penalty, the imposition of weak convexity here constrains its degree of nonconvexity. As previously mentioned, this ensures the proximal mapping of the penalty function remains single-valued, which is crucial for achieving superior algorithmic convergence. Combining the second order continuous differentiability of the loss and the assumption about the locally Lipschitz’s continuity of $p_{\lambda_{n}}^{{}^{\prime}}$ enables us to use the generalized Hessian to verify that a stationary point is a local minimizer. In fact, many concave penalties, including SCAD and MCP, possess locally Lipschitz continuous first derivatives on the positive half-line.
Moreover, in addition to the well-known SCAD and MCP regularizers, several other commonly used regularizers also satisfy the conditions in Assumption 1, we present a few examples below.
Example 1. Firm regularizer (Woodworth and Chartrand 2016):
$p_{\lambda_{n}}(t)=\left\{\begin{array}[]{rcl}\lambda_{n}|t|-\frac{\lambda_{n}t^{2}}{2c},&{|t|<c},\\ \frac{c\lambda_{n}}{2},&{|t|\geq c},\end{array}\right$.
where $\lambda_{n}$ is regularization parameter and $c>0$ is a fixed parameter. Assumption 1 holds with $\varrho=1$ and $\mu=\frac{1}{c}$.
@@
where $\lambda_{n}$ is regularization parameter and $c>0$ is a fixed parameter. Assumption 1 holds with $\varrho=\frac{1}{c}$ and $\mu=c$.
@@
where $\lambda_{n}$ is regularization parameter and $c>0$ is a fixed parameter. Assumption 1 holds with $\varrho=c$ and $\mu=c^{2}$.
@@
Let errors $\{\varepsilon_{1},\cdots,\varepsilon_{n}\}$ be independent and identically distributed Sub-Gaussian with variance proxy $\sigma^{2}$, and $\varepsilon_{i}$ have zero mean and positive variance $E\varepsilon_{i}^{2}$.
@@
For any $\bm{x},\bm{y}\in\mathbb{R}^{s}$, there exist constants $c_{4}\geq c_{3}>0$ satisfying that
@@
where $p_{1}=2\exp\left\{-s\ln(1+2n)\right\},p_{2}=\frac{1}{n^{-n/2}}$,
$r_{n}=C_{0}\left(\sqrt{\ln(1+2n)/n}+\lambda_{n}\varrho/\sqrt{s}\right)$ and $C_{0}=\max\left\{10\sigma\sqrt{c_{4}}/c_{1}c_{3},4/c_{1}^{2}c_{3}\right\}$.
The condition $c_{4}\geq 3/(ns)$ is non‑restrictive, as the right‑hand side diminishes to near zero for any practical $n$ and $s$, meaning any finite uniform bound $c_{4}$ will automatically satisfy it. Theorem 1 shows that the estimation error of the oracle estimator is bounded by $r_{n}$. Under the standard high‑dimensional scaling $(s\ln n+\ln d)/n\to 0$, one can choose $\lambda_{n}$ such that $\lambda_{n}/\sqrt{s}\to 0$, which ensures $r_{n}\to 0$ and thus the consistency of $\hat{\bm{\beta}}_{1}^{o}$ as $n\to\infty$.
@@
The inequalities in Assumption 4 are similar to the partial orthogonality condition proposed by Huang et al. (2008) for linear models, which was later extended to quadratic models by Fan et al. (2018). They can be regarded as that the sub-matrix of $\bm{Z}_{i}$ corresponding to $\Gamma^{\star}$ and the other complementary matrix being orthogonal.
@@
Under model (1) and Assumption 1-4, suppose
@@
then there exists a local minimizer $\hat{\bm{\beta}}$ of (2) such that
@@
    $\mathbb{P}(\hat{\bm{\beta}}_{{\Gamma^{\star}}^{c}}=0)\geq 1-p_{1}-2p_{2}-\frac{2}{n}$,
@@
The inequalities in (5) impose constraints on the scaling between $n$ and $s$. The first inequality, $36c_{2}^{2}c_{5}C_{0}s^{2}\leq\sqrt{n}$, implies $s=O(n^{1/4})$. The remaining two inequalities involve terms that decay to zero as $n$ grows and are thus satisfied for sufficiently large $n$.
@@
The condition $\mu\leq c_{1}^{2}c_{3}s$ restricts the weak convexity parameter $\mu$ of the penalty function, ensuring that it does not dominate the curvature of the loss function. For typical weakly convex penalties (e.g., SCAD, MCP), $\mu$ is a fixed constant. If $s\to\infty$, the right‑hand side grows unbounded, so a fixed $\mu$ satisfies the inequality for large $n$; if $s$ remains bounded, the condition imposes a mild fixed upper bound on $\mu$.

### v2

@@ 1 Introduction
Quadratic measurement regression model arises in numerous applications in physics, engineering, and data science, including phase retrieval (Candes et al. 2015), generalized phase retrieval (Wang and Xu 2019), the unassigned distance geometry problem (Huang and Dokmanić 2021), and power system state estimation (Wang et al. 2019). The essence is to recover the signals from noisy quadratic measurements data. Recent years have witnessed a surge of interests and methodological developments in this area. Notable advances include Thaker et al. (2020), Huang et al. (2020), Chen and Ng (2022), and Fan et al. (2025a), who proposed algorithms and theoretical guarantees of optimal solutions under various problem settings.
A particular exciting development is in the high-dimensional regime, where the dimension of the signal $d$ is large or even exceeds the number of sample points $n$. In this case, the sparsity assumption is crucial for ensuring identifiability and statistical efficiency. Several studies have explored this challenging direction, e.g., Bolte et al. (2018) studied the quadratic inverse problem and proposed a Bregman proximal gradient algorithm (PGA), while Zhang et al. (2023) and Ding et al. (2025) developed Bregman PGA for more general regularized problems. A special case of sparse quadratic model is sparse phase retrieval that attracted considerable attention (Soltanolkotabi 2019, Xia and Xu 2021, Cai et al. 2022, Huang and Xu 2024). More recently, Chen et al. (2025) employed the thresholded Wirtinger flow algorithm of Cai et al. (2016) for sparse phase retrieval.
However, most of the existing literature mentioned above focused primarily on numerical optimization and algorithmic development for the regularized least squares formulation of quadratic inverse problem, and the statistical guarantees remain relatively under-explored. On the other hand, in many applications the real data are noisy and in many cases, the noise level is significantly high relative to the signals. Therefore, investigation of statistical properties and stability of the methods and numerical procedures is desirable and important. A few researchers have attempted in this direction. Notably, Cai et al. (2016) developed a thresholded Wirtinger flow algorithm for noisy sparse phase retrieval that achieves a minimax optimal rate under sub-exponential noise satisfying a near-optimal sample size condition; while Fan et al. (2018) established a weak oracle property and proposed a fixed-point iterative algorithm for the $\ell_{q}(0<q<1)$-regularized least squares method.
In this paper, we consider the following general quadratic measurement model
@@
where $y_{i}\in\mathbb{R}$ denotes the observed response, $\bm{Z}_{i}\in\mathbb{R}^{d\times d}$ is a symmetric design matrix, $\bm{\beta}\in\mathbb{R}^{d}$ is the vector of unknown parameters representing true signals, and $\varepsilon_{i}\in\mathbb{R}$ is the random noise. Specifically, we study the regularization problem
@@
where $\lambda_{n}>0$ is a tuning parameter that encourages sparsity, and $P_{\lambda_{n}}(\cdot)$ is a penalty function. Model (1) is called generalized phase retrieval (Wang and Xu 2019, Chen and Ng 2022), whereas the phase retrieval is a special case where $\bm{Z}_{i}$ has rank one. This later case is also known as an index model in econometrics and statistics.
Various regularization techniques have been developed in the context of linear regression models, such as the LASSO, $\ell_{q}(0<q<1)$ penalty (Frank and Friedman 1993), SCAD (Fan and Li 2001), and MCP (Zhang 2010). It is well-known that LASSO has constant derivative that induces persistent shrinkage and biased estimation (Zou 2006). In contrast, concave penalties (e.g., SCAD, MCP, and $\ell_{q}(0<q<1)$) have derivatives that vanish for large coefficients, thereby mitigating shrinkage bias and yielding nearly unbiased estimates. However, nonconvex penalties introduces computational challenges. In particular, highly nonconvex penalties may cause numerous local minima, and the convergence of first-order algorithm (e.g., PGA) relies on favorable properties of the objective function. One such property is weak convexity, which is a key condition that generally ensures the proximal operator is single-valued under standard technical assumptions (Wang 2010, Khanh et al. 2025). If this condition is violated, the proximal operator may be multi-valued, which poses a potential risk that iterative algorithms could fail to converge to a meaningful solution.
This challenge motivates the study of regularizers that balance statistical and computational considerations. Loh and Wainwright (2015) developed a general framework for a class of weakly convex regularized M-estimators and established finite sample error bounds between any stationary point of the penalized estimator and the population-optimal solution without regularization (i.e., the solution minimizing the expected population risk). While their framework applies to a wide range of models, including linear and generalized linear models and graphical LASSO, it cannot be directly applied to sparse quadratic measurement problems. The main difficulty is that the least squares loss for this model leads to a highly nonconvex quartic polynomial that are computationally very challenging to minimize. In fact, as pointed out by Candes et al. (2015), it is already known that determining whether a stationary point of a quartic polynomial is a local minimizer is NP-hard. Moreover, Loh and Wainwright (2015) introduced an extra tuning parameter that must be chosen carefully to ensure that the true value $\bm{\beta}^{*}$ is a feasible point, thereby further increasing the difficulty of solving the problem. Weakly convex regularization methods have also been widely used in sparse signal recovery (Yang et al. 2019, Komuro et al. 2022) and linear inverse problems (Shumaylov et al. 2024, Goujon et al. 2024, Ebner et al. 2025), but these works do not explicitly exploit concave penalties.
To overcome the above mentioned numerical issues, in this paper, we propose a class of concave penalties that also satisfy weak convexity, called Weakly-Convex-Concave Penalties (WCCP). The major advantage of this class is that it provides a flexible framework for analyzing local minimizers in sparse quadratic measurement problems, combining the statistical benefits of concavity with the algorithmic guarantees of weak convexity.
Our contributions are threefold. First, we bridge a notable gap in the literature by providing the first systematic statistical analysis of the WCCP-regularized estimator for the quadratic measurement model (1). We establish rigorous support recovery guarantees and $\ell_{2}$-error bounds for its local minimizer, and derive the statistical properties of the proposed estimator. In particular, by choosing the appropriate regularization parameter, the $\ell_{2}$-error bounds achieve the rate of $O(\sqrt{\ln d/n})$, which is the known optimal rate for sparse phase retrieval as shown in Cai et al. (2016). Second, we develop iterative algorithms tailored to different regularization functions that provide convergence guarantees. Third, we demonstrate the effectiveness of our method, particularly in the context of sparse phase retrieval, and show that it performs comparable or better than the existing LASSO-based approaches.
@@ 2 Finite Sample Statistical Results
In this section, we derive the finite sample properties and consistency of the weakly convex-concave estimator in the high-dimensional case where $d$ is large than $n$. We assume that $\ln d=o(n^{\alpha})$ with some constant $\alpha\in(0,1)$ and denote the number of non-zero elements of the true signal by $s$. Following the literature (Huang et al. 2008, Fan et al. 2018), we assume that there exists constants $0<c_{1}\leq c_{2}<\infty$ such that
@@
Further assume that the observed response and design matrix are standardized so that
@@
The regularization function is coordinate separable
@@
for some scalar function $p_{\lambda_{n}}$ which satisfies:
@@
    $p_{\lambda_{n}}$ is concave on $(0,\infty)$ and there exists $\mu>0$ such that $p_{\lambda_{n}}(t)+\frac{\mu}{2}t^{2}$ is convex.
@@
    $p_{\lambda_{n}}$ satisfies $p_{\lambda_{n}}(0)=0$ and $p_{\lambda_{n}}(t)>0$ for $t>0$.
@@
    For $t>0$, $p_{\lambda_{n}}$ is non-decreasing and $t\mapsto\frac{p_{\lambda_{n}}(t)}{t}$ is non-increasing in $t$.
@@
    $p_{\lambda_{n}}$ is differentiable with derivative $p^{\prime}_{\lambda_{n}}$ for all $t>0$, and subdifferentiable at $t=0$, satisfying $\lim_{t\to 0^{+}}p^{\prime}_{\lambda_{n}}(t)=\lambda_{n}\varrho$ with $\varrho>0$. The function $p^{\prime}_{\lambda_{n}}$ is locally Lipschitz continuous in $(0,\infty)$.
Compared to Loh and Wainwright (2017), Loh (2017), here we requires the penalty function to be concave and its derivative $p_{\lambda_{n}}^{\prime}$ to be locally Lipschitz continuous on the positive half-line. We explicitly emphasize the concavity of the penalty function because it is known to possess stronger variable selection capabilities. However, we impose weak convexity here to ensure the proximal mapping of the penalty function remains single-valued, which is crucial for achieving superior algorithmic convergence. Combining the second order continuous differentiability of the loss and the local Lipschitz’s continuity of $p^{\prime}_{\lambda_{n}}$ enables us to use the generalized Hessian to verify that a stationary point is a local minimizer. In fact, many concave penalties, including SCAD and MCP, possess locally Lipschitz continuous first derivatives on the positive half-line. In the following we present some examples of regularizers satisfying Assumption 1, including SCAD, and some other commonly used regularizers.
Example 1. The Transformed $\ell_{1}$ (TL1) penalty (Lv and Fan 2009) is defined as
$p_{\lambda_{n}}(t)=\lambda_{n}\cdot\frac{(a+1)t}{a+t},\qquad t>0,$
where $a>0$ is a shape parameter. Assumption 1 holds with $\varrho=\dfrac{a+1}{a}$ and $\mu=2\lambda_{n}\frac{a+1}{a^{2}}$.
@@
where $\lambda_{n}$ is the regularization parameter and $c>0$ is a fixed parameter. Assumption 1 holds with $\varrho=c$ and $\mu=\lambda_{n}c^{2}$.
@@
where $\lambda_{n}$ is the regularization parameter and $c>0$ is a fixed parameter. Assumption 1 holds with $\varrho=c$ and $\mu=\lambda_{n}c^{2}$.
Example 4. The SCAD penalty is defined as
$p_{\lambda}(t)=\begin{cases}\lambda|t|,&|t|\leq\lambda,\\[4.0pt] -\dfrac{t^{2}-2a\lambda|t|+\lambda^{2}}{2(a-1)},&\lambda<|t|\leq a\lambda,\\[4.0pt] \dfrac{(a+1)\lambda^{2}}{2},&|t|>a\lambda,\end{cases}$
where $\lambda>0$ is the regularization parameter and $a>2$ is a fixed shape parameter. Assumption 1 holds with $\varrho=1$ and $\mu=1/(a-1)$.
Example 5. The MCP penalty takes the form
$p_{\lambda}(t)=\begin{cases}\lambda|t|-\dfrac{t^{2}}{2b},&|t|\leq b\lambda,\\[6.0pt] \dfrac{b\lambda^{2}}{2},&|t|>b\lambda,\end{cases}$
where $\lambda>0$ and $b>0$ is a shape parameter controlling the concavity. Assumption 1 holds with $\varrho=1$ and $\mu=1/b$.
The weak convexity parameters for SCAD and MCP shown in Examples 4 and 5 were previously established in Loh and Wainwright (2017) and Loh (2017).
@@
The errors $\{\varepsilon_{1},\cdots,\varepsilon_{n}\}$ are independent and identically distributed sub-Gaussian with variance proxy $\sigma^{2}$, and each $\varepsilon_{i}$ has zero mean and positive variance.
@@
For any $\bm{x},\bm{y}\in\mathbb{R}^{s}$, there exist constants $c_{4}\geq c_{3}>0$ satisfying
@@
where $p_{1}=2\exp\left\{-s\ln(1+2n)\right\},p_{2}=\frac{1}{n^{-n/2}}$, $r_{n}=C_{0}\left(\sqrt{\ln(1+2n)/n}+\lambda_{n}\varrho/\sqrt{s}\right)$ and $C_{0}=\max\left\{10\sigma\sqrt{c_{4}}/c_{1}c_{3},4/c_{1}^{2}c_{3}\right\}$.
The condition $c_{4}\geq 3/(ns)$ is not restrictive, since the right‑hand side diminishes to near zero for any practical $n$ and $s$, so that any finite uniform bound $c_{4}$ will automatically satisfy it. Theorem 1 shows that the estimation error of the oracle estimator is bounded by $r_{n}$. Under the standard high‑dimensional scaling $(s\ln n+\ln d)/n\to 0$, one can choose $\lambda_{n}$ such that $\lambda_{n}/\sqrt{s}\to 0$ to imply $r_{n}\to 0$ and thus the consistency of $\hat{\bm{\beta}}_{1}^{o}$ as $n\to\infty$.
@@
The inequalities in Assumption 4 are similar to the partial orthogonality condition in Huang et al. (2008) for linear models, which was later extended to quadratic models by Fan et al. (2018). They can be regarded as that the sub-matrix of $\bm{Z}_{i}$ corresponding to $\Gamma^{\star}$ and the complementary matrix being orthogonal.
@@
Under model (1) and Assumption 1-4, suppose that the triple $(n,s,d)$ satisfy
@@
there exists a local minimizer $\hat{\bm{\beta}}$ of (2) such that
@@
    $\mathbb{P}(\hat{\bm{\beta}}_{{\Gamma^{\star}}^{c}}=0)\geq 1-p_{1}-2p_{2}-\frac{2}{n}$; and
@@
The inequalities in (5) impose constraints on the scaling between $n$ and $s$. The first inequality, $36c_{2}^{2}c_{5}C_{0}s^{2}\leq\sqrt{n}$, implies $s=O(n^{1/4})$. The remaining two inequalities involve terms that decay to zero as $n$ grows and thus are satisfied for sufficiently large $n$.
@@
The condition $\mu\leq c_{1}^{2}c_{3}s$ restricts the weak convexity parameter $\mu$ of the penalty function, ensuring that it does not dominate the curvature of the loss function. For typical weakly convex penalties (e.g., SCAD, MCP), $\mu$ is a fixed constant. If $s\to\infty$, the right‑hand side grows unbounded, so a fixed $\mu$ satisfies the inequality for large $n$; if $s$ remains bounded, then the condition imposes a mild fixed upper bound on $\mu$.